OpenStudy (anonymous):

Two forces with magnitudes of 100 and 50 pounds act on an object at angles of 50° and 160° respectively. Find the direction and magnitude of these forces. Round to two decimal places in all intermediate steps and in your final answer.

OpenStudy (whpalmer4):

Draw yourself a careful diagram. Then use \(x = r \cos \theta\) and \(y = r \sin \theta\) to break them into their horizontal and vertical components. As a check, from the diagram you can see that the 100 lb force will be split about evenly going to +y and +x, and the 50 lb force will be mostly going to -x and a bit of +y.